# Marathon: A Race Through the Realm of Long Context with Large Language Models

### v1

@@ Abstract
Although there are currently many benchmarks available for evaluating the long context understanding and reasoning capability of large language models, with the expansion of the context window in these models, the existing long context benchmarks are no longer sufficient for evaluating the long context understanding and reasoning capability of large language models. In this paper, we have developed a fresh long context evaluation benchmark, which we name it Marathon in the form of multiple choice questions, inspired by benchmarks such as MMLU, for assessing the long context comprehension capability of large language models quickly, accurately, and objectively. We have evaluated several of the latest and most popular large language models, as well as three recent and effective long context optimization methods, on our benchmark. This showcases the long context reasoning and comprehension capabilities of these large language models and validates the effectiveness of these optimization methods. Marathon is available at https://huggingface.co/datasets/Lemoncoke/Marathon. [^1] We also provide an online evaluation website: https://openbenchmark.online/marathon
@@ 1 Introduction
Figure 1: An example of a test case in benchmark. The context is truncated for display purposes.
With the rapid development of artificial intelligence technology, large language models represented by GPT-4 (OpenAI, 2023b) have demonstrated powerful capabilities. In addition to the well-known ChatGPT, the application of large language models has gradually extended to various industries. The large model in the medical field, MedGPT, can help people consult medical-related questions, while the large model in the legal field, LawGPT (Nguyen, 2023), can assist people in legal consultations, significantly reducing the barriers and costs for people to obtain knowledge. The large model in the programming field, Github Copilot (Chen et al., 2021), can help developers complete their work faster, greatly improving productivity.
However, current large language models can only provide answers to users’ questions and requests within a relatively small context window. There are two main reasons for this limitation. Firstly, large language models still lack the ability to extract crucial information from long contexts. As described in Liu et al. (2023) large language models can only effectively retrieve key information when it is located at the beginning or end of the prompt. When the key information is in the middle of the prompt, it becomes difficult to retrieve successfully. This phenomenon becomes more evident as the context length increases. Secondly, longer contexts require more computational resources and time for large language models to provide feedback, which not only increases the hardware requirements but also reduces the user experience. Therefore, the ability of large language models to handle long contexts becomes increasingly important, and comprehensive evaluation of their performance in this regard is crucial. Existing long context benchmarks, such as Longbench (Bai et al., 2023b), still employ the F1 score as the evaluation metric, which measures the similarity between the generated response of the large language model and the candidate answers. However, the generated content of large language models is highly diverse, and the candidate answers cannot cover all possible correct responses, making the evaluation of large language models less accurate.
To address these challenges, we propose a new long context benchmark called Marathon based on Loogle (Li et al., 2023b), which includes 6 evaluation tasks for assessing the ability of large language models. The context lengths in the benchmark range from 60k to 260k+. For each long context, there is a related question and four carefully crafted answer options. The large language model needs to provide the correct answer option based on the given long context.
We conducted evaluations on Marathon using various large language models with different parameter sizes and context windows, combined with different methods of compressing long contexts. We found that the long context processing capabilities of most large language models still need improvement. In fact, the 7B Mistral (Jiang et al., 2023a) model outperformed the 70B Beluga (Mahan et al., 2023) model in terms of long context understanding. Some compression methods for long contexts were found to be effective in improving large language models’ understanding of long contexts in certain cases.
@@ 2 Related Work
2.1 Flash Attention
The training and inference of ultra-long sequences have been hindering the further development of large language models. In traditional Attention mechanisms, the GPU memory consumption of Attention Scores is proportional to the square of the sequence length. Therefore, when the sequence length reaches a certain threshold, both training and inference will encounter out-of-memory (OOM) issues. The emergence of Flash Attention (Dao et al., 2022) allows us to avoid computing the Attention Scores matrix and directly compute the final Attention Output result, greatly reducing the GPU memory required during model training and inference. Moreover, thanks to the reduction in the number of memory accesses between HMB (high-bandwidth memory) and GPU SRAM (static random-access memory) brought by Flash Attention, the speed of model training and inference is accelerated. Flash Attention 2 (Dao, 2023) further optimizes the implementation of Flash Attention, further accelerating the speed of model computation. Currently, Flash Attention 2 (Dao, 2023) has become a key technology for optimizing the training and inference process of large language models, playing a crucial role in modeling and inferring ultra-long sequences.
2.2 Long Context Models
The ability of large language models for handling long contexts has become increasingly important. ChatGPT (OpenAI, 2023a) supports a window size of 16k, while GPT-4 supports a window size of 128k, and Claude-2.1 supports a window size of 200k [^2] https://www.anthropic.com/index/claude-2-1. Many open-source large language models have started to expand the size of their context window. Longchat (Li et al., 2023a) and MPT (Team, 2023) have achieved a window size of 16k, while Mistral (Jiang et al., 2023a) and Zephyr (Tunstall et al., 2023) have achieved a window size of 32k. By utilizing an adapted Rotary Embedding (Su et al., 2022) and sliding window (Beltagy et al., 2020) during fine-tuning, MistralLite, based on Mistral, has achieved a window size of 128k, enabling large language models to handle even longer contextual information.
@@ 2.4 Retrieval Augmented Generation
Retrieval Augmented Generation (RAG) was originally proposed and applied to NLP tasks in (Lewis et al., 2020), and it has now become a mainstream method for improving the generation capability of large language models. RAG can extract the most relevant data from external knowledge bases and hand it over to the large language model for processing. This can alleviate the hallucination problem of large language models and enable people to trace the source of the content generated by large language models, ensuring the reliability of the generated content. Additionally, RAG can also be used to extract information from long documents that is most relevant to the user’s query. This ensures that key information required to provide correct answers to questions is not lost while reducing the length of the context. Many projects such as longchain (longchain, 2022) and LlamaIndex (Liu, 2022) have achieved significant progress in combining RAG with large language models, greatly facilitating related research in this direction.
@@ 2.5 Long Context Benchmarks
There have been many recent benchmarks used to assess the long context processing ability of large language models, such as Loogle (Li et al., 2023b) and LongBench (Bai et al., 2023b). Liu et al. (2023) on the other hand, noticed that the position of key information in long contexts greatly affects the capability of large language models to correctly understand and process text. Therefore, they used the NaturalQA (Kwiatkowski et al., 2019) dataset to construct a new benchmark to test the impact of different positions of key information in long context on the text processing capability of large language models.
Although Loogle (Li et al., 2023b) and LongBench (Bai et al., 2023b) have constructed a relatively comprehensive set of evaluation tasks, the evaluation metrics used are still F1-score, Bleu or Rouge, which cannot accurately evaluate the ability of large language models to handle and understand long contexts.
@@
Figure 2: The distribution of context length in Marathon.
3 Marathon
Currently, the mainstream benchmarks for evaluating large language models mostly adopt a multiple-choice question format, such as MMLU (Hendrycks et al., 2021) and C-Eval (Huang et al., 2023). The use of multiple-choice questions can effectively avoid scenarios where large language models generate correct answers but receive low scores due to the lack of relevant representations in the reference answers, or generate correct answers but receive high scores due to their similarity with the content of the reference answers. Therefore, based on Loogle (Li et al., 2023b) and Longbench (Bai et al., 2023b), we have constructed a multiple-choice long-text benchmark to more reasonably assess the long-context comprehension ability of large language models.
3.1 Overview
The Marathon benchmark consists of 6 tasks: Comprehension and Reasoning, Multiple Information Retrieval, Timeline Reorder, Computation, Passage Retrieval, and Short Dependency QA. These tasks can be categorized into four types based on the nature of the questions: Question Answering, Timeline Reorder, Computation, and Passage Retrieval. Figure 1 shows sample test cases for each type of question. Table 1 shows the quantity information of test samples for each task.
@@ 3.2 Construction
All the test samples in the benchmark are in the form of multiple-choice questions, with each question containing one correct answer option and several distractor options. We use GPT-4 to generate the distractor options for each question. For each question, we divide the long context into multiple fragments of length 12,000 and randomly select one fragment. We require GPT-4 to generate three distractor options based on the given context fragment, question, and correct answer. The purpose of this approach is to avoid using excessively long context that exceeds GPT-4’s context window, which may affect the accuracy of the generated results. By using shorter contexts, we can obtain distractor options that are more relevant to these shorter contexts. Finally, to ensure the effectiveness and accuracy of these distractor options, we manually verify the options of each test sample.
@@ 3.3 Question Answering
Comprehension and Reasoning, Multiple Information Retrieval and Short Dependency QA are all types of traditional question-answer formats. The difference lies in the fact that Comprehension and Reasoning, Multiple Information Retrieval are selected from the Long Dependency QA dataset in Loogle (Li et al., 2023b), while Short Dependency QA is selected from the Short Dependency QA dataset in Loogle (Li et al., 2023b). In the question answering tasks, each question is accompanied by a corresponding long context, and the large language model is required to infer the correct answer according to the long context. For the Short Dependency QA task, the relevant content for the correct answer is relatively concentrated within the long context. For Comprehension and Reasoning and Multiple Information Retrieval tasks, the content relevant to the correct answer is more scattered throughout the long context. Therefore, the large language model needs to possess strong long context understanding capability in order to solve the question correctly.
@@ 3.4 Timeline Reorder
Timeline Reorder task is a relatively novel question answering task. Unlike traditional question answering tasks, in the Timeline Reorder task, the question format requires large language models to sort a series of events described in a long context according to their chronological order. This task aims to examine the large language models’ understanding of temporal relationships. Due to the dispersed distribution of events that need to be sorted by chronological order in the long context, large language models not only need to possess a correct understanding of temporal order but also require strong long context processing capabilities to answer correctly, which makes it a challenging task.
@@ 3.5 Computation
Computation task is also different from traditional question answering tasks. Its question format involves providing a question related to numerical computation and requires the large language model to perform numerical calculations based on relevant content in the long context. For example, it may require calculating the number of children a certain character has at a specific time point, considering that the long context describes the character’s life events, including the death of a child due to illness, which may affect the number of the character’s offspring at subsequent time points. Therefore, to answer this question correctly, the large language model not only needs to be able to perform ordinary numerical calculations but also needs to capture all the key information related to the question. Compared to traditional computation and question answering tasks, this task is more challenging and can better reflect the large language model’s capability to comprehend long context.

### v2

@@ Abstract
With the advancement of large language models (LLMs) and the expansion of their context windows, existing long-context benchmarks fall short in effectively evaluating the models’ comprehension and reasoning abilities in extended texts. Moreover, conventional benchmarks relying on F1 metrics often inaccurately score responses: they may undervalue correct answers that differ from the reference responses and overvalue incorrect ones that resemble the reference texts. In response to these limitations, we introduce Marathon, a novel evaluation benchmark that adopts a multiple-choice question format. It is specifically designed to overcome the constraints of previous benchmarks and provide a rapid, precise, and unbiased appraisal of the long-context comprehension skills of large language models. We conducted comprehensive evaluations on the Marathon benchmark with a range of state-of-the-art LLMs and assessed the effectiveness of various optimization strategies tailored for long-context generation. We anticipate that the Marathon benchmark and its associated leaderboard will enable a more precise and equitable evaluation of LLMs’ capabilities in understanding and reasoning over extended contexts. Marathon is available at https://github.com/Hambaobao/Marathon. [^1] We also provide an online evaluation website: https://openbenchmark.online/marathon
@@ 1 Introduction
Figure 1: The overall accuracy of different models on Marathon. The x-axis represents the model, and the y-axis represents the average accuracy across all tasks. The different colors represent different methods of optimization.
In the rapidly evolving landscape of artificial intelligence technologies, the emergence of large language models (LLMs), as exemplified by ChatGPT (OpenAI et al., 2024), showcases notable capabilities. The influence of these models extends beyond the well-established ChatGPT, gaining increasing prominence across diverse sectors. Existing LLMs are typically built upon Transformer architectures, which demand memory and computational resources that grow quadratically with sequence length. Consequently, Transformer language models have historically been trained with relatively modest predetermined context windows. For instance, LLaMA Touvron et al. (2023a) employs a context size of 2048 tokens, while Llama2 Touvron et al. (2023b) utilizes a context size of 4096 tokens. However, the pre-defined size imposes constraints on LLMs in various applications, such as summarizing extensive documents or addressing lengthy questions.
Significant research efforts have been devoted to extending the context length of LLMs. Due to the prohibitive expense of training LLMs with extended context lengths from scratch, the predominant studies have endeavored to enhance the capabilities of LLMs to comprehend long contexts through fine-tuning. These methods encompass extending the context window Chen et al. (2023b), incorporating recurrent memory Bulatov et al. (2024), employing sparse attention mechanisms Xiao et al. (2023a), and augmenting with external memory Wang et al. (2023). Concurrently, an increasing multitude of benchmarks have been introduced to assess the long-context understanding capabilities of LLMs. LongBench Bai et al. (2023b) stands out as the first bilingual, multi-task benchmark specifically designed for the assessment of long-context understanding. This dataset continues to depend on the F1 score, which evaluates the responses of LLMs against a predefined set of possible answers. LooGLE Li et al. (2023b) encompasses intricate long dependency tasks, including event timeline reordering, comprehension/reasoning, and computation. Nevertheless, the diverse nature of model-generated content introduces a challenge, as these predefined answers may not encompass all valid responses, thereby diminishing the precision of assessing model performance. There is a growing demand for high-quality benchmarks characterized by significantly longer text lengths and more challenging tasks, ensuring comprehensive evaluations.
In this study, we introduce a novel benchmark named Marathon, designed for long-context understanding and reasoning. In particular, this benchmark is constructed upon the foundations established by LooGLE (Li et al., 2023b) and LongBench (Bai et al., 2023b). The contextual lengths within this benchmark span from 2K to over 260K characters. For each extensive context provided, an associated question is paired with four meticulously crafted response options. These options have been carefully reviewed by humans and contain only one correct answer, with the remaining options designed to be highly misleading. This design makes the Marathon benchmark a particularly challenging one. The task for the large language model is to discern the accurate response option based on the extensive context provided.
The main contributions of this work are threefold:
  • We introduce a novel multiple-choice long context benchmark that comprehensively evaluates the long context understanding and reasoning capabilities across 10 leading open-source large language models, as well as ChatGPT and GPT-4, covering six diverse types of tasks.
  • We compare two prevalent methods for long context optimization (Prompt Compression and Retrieval Augmented Generation) along with two leading embedding models, assessing their impact on enhancing the long context reasoning abilities of large language models.
  • Our findings reveal a general tendency among current open-source large language models to generate lengthier responses, accompanied by a notable deficiency in following instructions accurately.
@@ 2.2 Retrieval Augmented Generation
Retrieval Augmented Generation (RAG) was originally proposed and applied to NLP tasks in (Lewis et al., 2020), and it has now become a mainstream method for improving the generation capability of large language models. RAG can extract the most relevant data from external knowledge bases and hand it over to the large language model for processing. This can alleviate the hallucination problem of large language models and enable people to trace the source of the content generated by large language models, ensuring the reliability of the generated content. Additionally, RAG can also be used to extract information from long documents that is most relevant to the user’s query. This ensures that key information required to provide correct answers to questions is not lost while reducing the length of the context. Many projects such as Longchain  [^2] https://github.com/langchain-ai/langchain and LlamaIndex  [^3] https://github.com/jerryjliu/llama_index have achieved significant progress in combining RAG with large language models, greatly facilitating related research in this direction.
2.3 Long Context Models
The ability of large language models for handling long contexts has become increasingly important. ChatGPT  [^4] https://chat.openai.com supports a window size of 16k, while GPT-4 supports a window size of 128k, and Claude-2.1 supports a window size of 200k [^5] https://www.anthropic.com/index/claude-2-1. Many open-source large language models have started to expand the size of their context window. Longchat (Li et al., 2023a) and MPT (Team, 2023b) have achieved a window size of 16k, while Mistral (Jiang et al., 2023a) and Zephyr (Tunstall et al., 2023) have achieved a window size of 32k. By utilizing an adapted Rotary Embedding (Su et al., 2022) and sliding window (Beltagy et al., 2020) during fine-tuning, MistralLite, based on Mistral, has achieved a window size of 128k, enabling large language models to handle even longer contextual information.
Figure 2: Examples of test cases in the benchmark. The context is truncated for display purposes.
@@ 2.4 Long Context Benchmarks
There have been many recent benchmarks (Tay et al., 2020; Pang et al., 2022; Shaham et al., 2022) used to assess the long context processing ability of large language models, such as LooGLE (Li et al., 2023b) and LongBench (Bai et al., 2023b). Liu et al. (2023b) on the other hand, noticed that the position of key information in long contexts greatly affects the capability of large language models to correctly understand and process text. Therefore, they used the NaturalQA (Kwiatkowski et al., 2019) dataset to construct a new benchmark to test the impact of different positions of key information in long context on the text processing capability of large language models.
Pang et al. (2022) proposed a multiple-choice long context benchmark, but its average text length is only 5k, which is significantly below the context window of mainstream large language models. Although SCROLLS (Shaham et al., 2022), LooGLE (Li et al., 2023b), and LongBench (Bai et al., 2023b) have constructed a relatively comprehensive set of evaluation tasks, the evaluation metrics used are still F1-score, Bleu or Rouge, which cannot accurately evaluate the ability of large language models to handle and understand long contexts.
3 Marathon
Present benchmarks for evaluating large language models primarily use a multiple-choice format, highlighted by studies such as MMLU (Hendrycks et al., 2021) and C-Eval (Huang et al., 2023). This multiple-choice approach helps prevent situations where large language models produce correct answers but are scored low due to missing corresponding elements in the reference answers or when they produce incorrect answers that are scored high because they resemble the reference answers closely. Therefore, influenced by LooGLE (Li et al., 2023b) and LongBench (Bai et al., 2023b), we developed a multiple-choice, long-context benchmark to more accurately evaluate the ability of large language models to understand extended contexts.
3.1 Overview
The Marathon benchmark includes six tasks: Comprehension and Reasoning, Multiple Information Retrieval, Timeline Reorder, Computation, Passage Retrieval, and Short Dependency QA. These tasks are grouped into four categories based on the type of questions they involve: Question Answering, Timeline Reordering, Computation, and Passage Retrieval. Table 1 provides the number of test samples for each task. Figure 2 presents example questions for each category.
@@ 3.2 Construction
All the test samples in the benchmark are in the form of multiple-choice questions, with each question containing one correct answer option and several distractor options. We use GPT-4 to generate the distractor options for each question. For each question, we divide the long context into multiple fragments of length 12,000 tokens and randomly select one fragment. We require GPT-4 to generate three distractor options based on the given context fragment, question, and correct answer. The purpose of this approach is to avoid using excessively long context that exceeds GPT-4’s context window, which may affect the accuracy of the generated results. By using shorter contexts, we can obtain distractor options that are more relevant to these shorter contexts. The distractor options for all test samples were generated using the OpenAI GPT-4-0613 API, with a total expenditure of approximately \$900.
Finally, to ensure the effectiveness and accuracy of these distractor options, we manually verify the options of each test sample.
@@ 3.3 Question Answering
Comprehension and Reasoning, Multiple Information Retrieval and Short Dependency QA are all types of traditional question-answer formats. The difference lies in the fact that Comprehension and Reasoning, Multiple Information Retrieval are selected from the Long Dependency QA dataset in LooGLE (Li et al., 2023b), while Short Dependency QA is selected from the Short Dependency QA dataset in LooGLE (Li et al., 2023b). In the question-answering tasks, each question is accompanied by a corresponding long context, and the large language model is required to infer the correct answer according to the long context. For the Short Dependency QA task, the relevant content for the correct answer is relatively concentrated within the long context. For Comprehension and Reasoning and Multiple Information Retrieval tasks, the content relevant to the correct answer is more scattered throughout the long context. Therefore, the large language model needs to possess strong long context understanding capability in order to solve the question correctly.
@@
Figure 3: The distribution of context lengths for six tasks in the Marathon benchmark.
@@ 3.4 Timeline Reorder
Timeline Reorder task is a relatively novel question-answering task. Unlike traditional question-answering tasks, in the Timeline Reorder task, the question format requires large language models to sort a series of events described in a long context according to their chronological order. This task aims to examine the large language models’ understanding of temporal relationships. Due to the dispersed distribution of events that need to be sorted by chronological order in the long context, large language models not only need to possess a correct understanding of temporal order but also require strong long context processing capabilities to answer correctly, which makes it a challenging task.
@@ 3.5 Computation
Computation task is also different from traditional question-answering tasks. Its question format involves providing a question related to numerical computation and requires the large language model to perform numerical calculations based on relevant content in the long context. For example, it may require calculating the number of children a certain character has at a specific time point, considering that the long context describes the character’s life events, including the death of a child due to illness, which may affect the number of the character’s offspring at subsequent time points. Therefore, to answer this question correctly, the large language model not only needs to be able to perform ordinary numerical calculations but also needs to capture all the key information related to the question. Compared to traditional computation and question answering tasks, this task is more challenging and can better reflect the large language model’s capability to comprehend long context.